OpenStudy (anonymous):

If the value of sin θ= -0.798, which of the following could be true about θ? 0 ≤ θ ≤ π/3 π/3 ≤ θ < 2π/3 2π/3 ≤ θ < 5π/4 5π/4 ≤ θ < 7π/4 7π/4 ≤ θ < 2π

OpenStudy (campbell_st):

look at where sin is negative 3rd quadrant \[(\pi + \theta)\] and 4th quadrants... \[(2\pi - \theta)\] find the radian value of \[\sin \theta = 0.798\] just identify which inequalities

OpenStudy (anonymous):

i still dont understand, can u break it down any further, plz?

OpenStudy (campbell_st):

what is sin (theta) = 0.798

OpenStudy (campbell_st):

nope... find \[\theta = \sin ^{1}(0.798\]

OpenStudy (campbell_st):

do you know about the unit circle and where trig ratios are positive and negative..?